A force of 490 newton acts on a 5.0 kilogram mass for 12 second s to bring it to rest. Find that acceleration produced
1
Expert's answer
2020-12-03T10:25:23-0500

We can find the acceleration produced by the force from the Newton's Second Law of Motion:


"F=ma,""a=\dfrac{F}{m}=\dfrac{490\ N}{5.0\ kg}=98\ \dfrac{m}{s^2}."

Answer:

"a=98\ \dfrac{m}{s^2}."